Match up the structures with the Spectrums, (A-E), explain why you matched the structure with the spectrum (ex: a large broad signal at approximately 3500cm^-1 so you have chosen an alcohol), and give the name of the IUPAC name of the compound.Structures: OH OH

Spectrum A: IOD TPANSTANCIU มา 4000 9000 2000 1500 1000 HAVE NUMBERSpectrum B: 1000 3000 2000 Wavenumber (cm-1)Spectrum C: 3000 1000 2000 Wavenumber (cm-1)Spectrum D: 3000 2000 Wavenumber (cm-1) 1000Spectrum E: 0 . 8 ا ليبياه.0 3000 1000 2000 Wavenumber (cm-1)
Answer #1

Hi five compounds are different, so we can assign them unambiguously.

IUPAC names are provided as

Spectrm-X: IUPAC name

---------------------------------------------------------------------------------------------------------------

Spectrum-A: butanal

1. strong peak in the region 1740-1720 cm-1 due to C=O stretch

2. medium two peaks in the region 2820-2850 & 2720-2750 cm-1 for =C-H stretch

---------------------------------------------------------------------------------------------------------------

Spectrum-B: Cyclopentanol

1. strong, broad signal in the region 3200-3600 cm-1 due to O-H stretch

2. strong peaks in the region 2850-3000 cm-1 due to sp3 alkyl C-H stretch

3. strong peak in the region 1050-1150 cm-1 due to C-O stretch

---------------------------------------------------------------------------------------------------------------

Spectrum-C: hexan-2-one

1. strong peaks in the region 2850-3000 cm-1 due to sp3 alkyl C-H stretch

2. strong peak in the region 1705-1725 cm-1 for C=O stretch

---------------------------------------------------------------------------------------------------------------

Spectrum-D: 1-Propoxypropane

1. strong peaks in the region 2850-3000 cm-1 due to sp3 alkyl C-H stretch

2. strong band in the region 1000-1300 (1070-1150) cm-1 due to C-O stretch

---------------------------------------------------------------------------------------------------------------

Spectrum-E: butanoic acid

1. strong peaks in the region 2850-3000 cm-1 due to sp3 alkyl C-H stretch

2. strong, very broad signal in the region 2500-3300 cm-1 due to O-H stretch

3. strong signal in the region 1700-1725 cm-1 due to C=O stretch

4. strong signal in the region 1210-1320 cm-1 due to C-O stretch

---------------------------------------------------------------------------------------------------------------

Spectrum 2000 2 000 Wavenumber (cm-1) 1000 Spectrum 3000 1000 2000 Wavenumber (cm-1) 3000 1000 2000 Wavenumber(cm-1) Spectrum

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")